OURO 4 wad of 10 R AN dy AYDAVE and den PN poupuran wing var en Nedrive 22 devene me enseny (U>100 kind CHIEDENIS 7:20
Answer #1

Given,

layer 1 : Cu = 100 Kn/m2

Y = 17 Kn/m2

L1 = 7m
For \phi = 0

Nc = 1.00

Nq = 5.70

Ny = 0.0

For circular footings: Sc = 1.3,

Sq = 1,

Sy = 0.6

diameter d = .6m

Bearing capacity = cNcSc + Y*L1*NqSq + .5Y*d*Ny*Sy

= 100 Kn/m2 x 1 x 1.3 +    17 Kn/m2 x 7m x 5.7 x 1   

= 808.3 Kn/m2

Layer 2: \phi = 26

Nc = 14.21

Nq = 27.06

Ny = 11.7

Y = 19 Kn/m2

L2 = 8m

c = 0

Bearing capacity = cNc + Y*L2*NqSq + .5Y*d*Ny*Sy

= 19 Kn/m2 x 8m x 27.06 x 1    +    .5 x 19 Kn/m2 x .6m x 11.7 x .6

= 4153.13 Kn/m2

Layer 3:

c =150 Kn/m2

Y = 20 Kn/m2

L3 = 7m

For \phi = 0

Nc = 1.00

Nq = 5.70

Ny = 0.0

Bearing capacity = cNcSc + Y*L3*NqSq + .5Y*d*Ny*Sy

= 150 Kn/m2 x 1 x 1.3    +    20 Kn/m2 x 7m x 5.7 x 1

= 993 Kn/m2

total bearing capacity pile footing = 808.3 Kn/m2 + 4153.13 Kn/m2 + 993 Kn/m2

= 5954.43 Kn/m2 for 1 pile

Now area of 1 footing = \pi /4 x d2

= .2827 m2

Load on 1 pile = bearing capacity x area

= 5954.43 Kn/m2 x .2827 m2

= 1683.57 KN

So, number of piles required for 20000 KN

= 20000/1683.57

= 11.87

so provide with 12 piles.

now total bearing capacity of pile group = 12 x 1683.57 KN

= 20202.84 KN

Efficiency = 200000/20202.8

= .989

= 98.9%